# On the Identifiability of Steering Vectors in Large Language Models

### v1

@@ 6.3 Scale Invariance Analysis
The curves track closely with overlapping confidence bands, demonstrating that $v$ and $v+v_{\perp}$ produce statistically indistinguishable behavioral effects at all tested steering strengths. For Qwen2.5-3B, mean differences remain below $0.022$ across all $\alpha$ values (less than $2.5\%$ deviation in formality scores); for Llama-3.1-8B, differences remain below $0.023$ ($2.5\%$ deviation). This confirms that non-identifiability is a structural property: any vector in the equivalence class $v+\ker(J)$ produces identical observations regardless of scaling.
@@ 7 Conclusion
Rather than viewing non-identifiability as a limitation, we frame it as a necessary foundation for principled steering research. Our characterization clarifies what behavioral observations can reveal about representations and provides a framework for designing interventions that are both empirically effective and theoretically grounded. The empirical validation demonstrates that these are not abstract concerns: non-identifiability affects real steering interventions in deployed models, with typical steering vectors containing substantial observationally irrelevant components. However, the constructive characterization in Proposition 2 shows that identifiability is achievable through principled design choices, enabling alignment methods that clearly specify their affordances and limitations.
@@ 8 Limitations
Our empirical validation covers two models at mid-network layers across three semantic traits (formality, politeness, humor). While our theoretical results apply generally, the empirical magnitude of non-identifiability, including null space dimensionality and the fraction of vector norm therein, may vary across model families, scales, architectures and layer positions. Our semantic evaluation uses lexical heuristics rather than full distributional metrics or human judgments, providing a conservative test of observational equivalence. The primary theoretical mechanism relies on local linear approximation, formally valid in a neighborhood of the reference distribution, though our empirical validation confirms large equivalence classes exist in practice.
Proposition 2 characterizes sufficient conditions for identifiability but comprehensive empirical evaluation of each regime across multiple models, tasks and experimental designs remains future work. The practical applicability of identifiable regimes involves trade-offs: ICA requires carefully curated contrastive prompts, sparsity regularization may reduce efficacy if true factors are not sparse and multi-environment validation requires substantial additional data. Understanding these trade-offs and extending our framework to other intervention modalities (activation patching, prompt based steering) are important directions for the representational alignment community.
@@ 9 Future Work
Systematic evaluation of identifiable regimes, comparing ICA-based extraction, sparse recovery, multi-environment training and cross-layer validation against standard contrastive methods, would clarify which structural assumptions are most effective in practice. Hybrid methods combining multiple identifiability conditions may achieve stronger identifiability than individual approaches, while layer-wise analysis could reveal whether certain network positions afford more identifiable steering and decomposing the Jacobian by architectural component could identify which elements contribute to non-identifiability.
The multi-environment identifiability condition connects to causal representation learning, where adapting invariant risk minimization to train steering vectors that maintain consistent effects across distributions could filter spurious correlations. Investigating adversarial robustness and out-of-distribution generalization would test whether identifiable methods exhibit superior robustness. Characterizing scaling laws for identifiability, including how null-space dimensionality evolves as models scale, would inform long-term strategy for controlling future systems.
Acknowledgments
@@
Appendix A Prompt Construction Details
A.1 Contrastive Prompt Pairs
We construct contrastive prompt pairs for each semantic trait using template-based generation. Figure 3 illustrates the general structure and Table 2 provides concrete examples.
Topic Pool 50 topics e.g., “weather”, “tech trends” Positive Template “formal”, “polite” Negative Template “informal”, “rude” 50 Contrastive Prompt Pairs
Figure 3: Contrastive prompt pair construction.
Trait Positive Prompt ($x^{+}$) Negative Prompt ($x^{-}$)
Formality Write a professional and formal message about technology trends. Write a casual and informal message about technology trends.
Politeness Write a polite and courteous response for a disagreement with someone. Write a rude and disrespectful response for a disagreement with someone.
Humor Write a humorous and funny response about an awkward moment. Write a serious and straightforward response about an awkward moment.
Table 2: Example contrastive prompt pairs for each semantic trait. Each pair consists of a positive prompt ($x^{+}$) designed to elicit high trait values and a negative prompt ($x^{-}$) designed to elicit low trait values.
A.2 Steering Extraction and Evaluation Pipeline
@@
Steering Extraction 50 contrastive pairs (trait-specific templates) Steering Vector $v\in\mathbb{R}^{d}$ Evaluation 100 held-out prompts (diverse, neutral) Steered Output $o(x,v,\alpha)$ extractapplygenerate Key: Test prompts are intentionally orthogonal to extraction prompts to ensure steering generalizes beyond training distribution.
Figure 4: Steering extraction and evaluation pipeline. Steering vectors are extracted from trait-specific contrastive pairs, then evaluated on diverse held-out prompts to test generalization.
Appendix B Detailed Proof of Proposition 1
B.1 Statement and Overview
Proposition 1.
Proof strategy.
We establish non-identifiability through two complementary mechanisms:
  • Null-space ambiguity (primary, constructive).
  • Reparameterization symmetry (existence-based).
B.2 Null-Space Ambiguity (Constructive Proof)
@@
$o(x,v^{\prime},\alpha)$ $\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}$ (9)
$=o(x,0,0)+\alpha J_{\ell}(x)v$
$\approx o(x,v,\alpha).$
@@
B.3 Reparameterization Symmetry (Existence Proof)
@@
$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)$ (16)
$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
$=h_{\ell+1}.$
@@
$\tilde{h}^{\prime}_{\ell}$ $=h^{\prime}_{\ell}+\alpha v^{\prime}$ (19)
$=Ah_{\ell}+\alpha Av$
$=A(h_{\ell}+\alpha v)$
$=A\tilde{h}_{\ell}$,
$\tilde{h}^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}(h^{\prime}_{\ell}+\alpha v^{\prime})+b^{\prime}_{\ell+1}\right)$
@@
Appendix C Detailed Proof of Proposition 2
C.1 Statement and Overview
Proposition 2.
Persona vectors can be identified up to scaling and permutation, thus affording reliable alignment control, under the following sufficient structural conditions:
  • Statistical Independence (ICA). Latent components are statistically independent, allowing unique recovery via independent component analysis.
  • Sparsity constraints. Steering directions admit sparse representations that reduce null-space ambiguity.
  • Multi-environment or interventional data. Variation across environments or interventions breaks observational equivalences.
  • Cross-layer consistency. Valid semantic directions propagate coherently across layers, filtering spurious components.
We provide detailed proofs for each condition.
C.2 Proof of Condition: Statistical Independence (ICA)
Setup: Assume the latent persona is a vector $z=(z_{1},\ldots,z_{k})\in\mathbb{R}^{k}$ with independent components and the representation follows:
$h_{\ell}=Az+\epsilon$ (20)
where $A\in\mathbb{R}^{d\times k}$ is a mixing matrix and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ is Gaussian noise.
Goal: Show that $A$ (and hence its columns, which are the steering vectors) can be recovered up to permutation and scaling.
Step 1: ICA identifiability theorem.
Let $x=As$ where $s\in\mathbb{R}^{k}$ has independent components and $A\in\mathbb{R}^{d\times k}$ is full column rank. Under the following sufficient conditions (Comon, 1994; Hyvärinen and Oja, 2000):
  • At most one component of $s$ is Gaussian
  • Components are statistically independent
  • Sufficient samples are observed
Then $A$ is identifiable up to column permutation and column scaling.
Important caveat: These are strong assumptions that may not hold exactly in practice for LLM personas. Statistical independence between semantic factors (e.g., formality and politeness) is an idealization; real personas may exhibit weak dependencies. The ICA framework provides a sufficient structural condition for identifiability, not a characterization of what typically holds in contemporary steering pipelines.
Step 2: Application to steering.
In our setting:
  • Observations: $h_{\ell}(x_{1}),\ldots,h_{\ell}(x_{N})$
  • Model: $h_{\ell}(x_{i})=Az(x_{i})+\epsilon_{i}$
  • Sources: $z(x_{1}),\ldots,z(x_{N})$ are realizations of independent persona components
Assumption verification:
  • Independence: We assume $z_{1},\ldots,z_{k}$ are statistically independent
  • Non-Gaussianity: At most one $z_{i}$ is Gaussian (e.g., politeness and formality are typically non-Gaussian in natural language)
  • Full rank: $A$ has full column rank (each persona has a non-zero effect on representations)
Step 3: Recovery via ICA algorithm.
Apply ICA algorithm (Hyvärinen and Oja, 2000) to observations $\{h_{\ell}(x_{i})\}_{i=1}^{N}$:
  • Whitening: Compute $\tilde{h}_{\ell}=\Sigma^{-1/2}(h_{\ell}-\mu)$ where $\mu=\mathbb{E}[h_{\ell}]$ and $\Sigma=\operatorname{Cov}(h_{\ell})$.
  • Independent component extraction: Find unmixing matrix $W$ that maximizes non-Gaussianity of $\hat{z}=W\tilde{h}_{\ell}$.
  • Recover mixing matrix: $\hat{A}=\Sigma^{1/2}W^{-1}$.
Step 4: Identifiability guarantee.
By the ICA theorem, $\hat{A}=APD$ where:
  • $P\in\mathbb{R}^{k\times k}$ is a permutation matrix
  • $D\in\mathbb{R}^{k\times k}$ is a diagonal scaling matrix
Therefore, the columns of $\hat{A}$ are the columns of $A$ up to permutation and scaling. Since steering vectors are defined as $v_{i}=Ae_{i}$ (the $i$-th column of $A$), we recover the true steering directions uniquely (up to unavoidable symmetries).
Step 5: Noise robustness.
With Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$, ICA remains consistent:
$h_{\ell}=Az+\epsilon$ (21)
Since $\epsilon$ is Gaussian and $z$ is non-Gaussian (by assumption), the ICA algorithm separates signal from noise:
  • Recovered sources: $\hat{z}=z+\tilde{\epsilon}$ where $\tilde{\epsilon}$ is small for $\sigma^{2}\ll\|Az\|^{2}$
  • Recovered mixing: $\hat{A}\approx APD$ with estimation error decreasing as $N\to\infty$
Conclusion (ICA): Under statistical independence, non-Gaussianity and full column rank—strong sufficient conditions that may not hold automatically in practice—persona vectors are identifiable up to permutation and scaling. These assumptions provide a theoretical pathway to identifiability but require careful experimental design to approximate in real steering settings.
C.3 Proof of Condition: Sparsity Constraints
Setup: Assume the true persona vector $v\in\mathbb{R}^{d}$ is sparse with $\|v\|_{0}=s\ll d$, meaning at most $s$ entries are non-zero. We observe the effect of steering:
$y=J_{\ell}v+\eta$ (22)
where $y\in\mathbb{R}^{m}$ are output measurements, $J_{\ell}\in\mathbb{R}^{m\times d}$ is the measurement matrix (Jacobian) and $\eta\sim\mathcal{N}(0,\sigma^{2}I)$ is noise.
Goal: Show that $v$ can be uniquely recovered via $\ell_{1}$ minimization.
Step 1: Compressed sensing setup.
The recovery problem is:
$\min_{v^{\prime}\in\mathbb{R}^{d}}\|v^{\prime}\|_{1}\quad\text{subject to}\quad\|J_{\ell}v^{\prime}-y\|_{2}\leq\epsilon$ (23)
where $\epsilon$ bounds the noise: $\epsilon\geq\|\eta\|_{2}$ with high probability.
Step 2: Restricted Isometry Property (RIP).
A matrix $J_{\ell}$ satisfies the RIP of order $s$ with constant $\delta_{s}$ if for all $s$-sparse vectors $v$:
$(1-\delta_{s})\|v\|_{2}^{2}\leq\|J_{\ell}v\|_{2}^{2}\leq(1+\delta_{s})\|v\|_{2}^{2}$ (24)
Theorem (Candes and Tao, 2005): If $J_{\ell}$ satisfies RIP with $\delta_{2s}<\sqrt{2}-1\approx 0.414$, then $\ell_{1}$ minimization recovers $v$ exactly (in the noiseless case) or approximately (with noise) with error:
$\|v-\hat{v}\|_{2}\leq C\epsilon$ (25)
for some constant $C$ depending on $\delta_{2s}$.
Important caveat: The RIP condition is a strong assumption. For Jacobians $J_{\ell}$ arising from neural network steering, RIP typically does not hold automatically and must be verified empirically or enforced through measurement design (diverse prompt selection). This condition is sufficient for identifiability but may not be satisfied in standard steering settings without careful experimental design.
Step 3: Recovery guarantee.
Solve:
$\hat{v}=\arg\min_{v^{\prime}}\|v^{\prime}\|_{1}\quad\text{s.t.}\quad\|J_{\ell}v^{\prime}-y\|_{2}\leq\epsilon$ (26)
By the compressed sensing theorem:
$\|v-\hat{v}\|_{2}\leq C_{1}\epsilon+C_{2}\frac{\|v-v_{s}\|_{1}}{\sqrt{s}}$ (27)
where $v_{s}$ is the best $s$-sparse approximation to $v$. If $v$ is exactly $s$-sparse, the second term vanishes and:
$\|v-\hat{v}\|_{2}\leq C_{1}\epsilon$ (28)
Step 4: Uniqueness.
Suppose two sparse vectors $v$ and $v^{\prime}$ both satisfy $\|J_{\ell}v-y\|_{2}\leq\epsilon$. Then:
$\|J_{\ell}(v-v^{\prime})\|_{2}\leq 2\epsilon$ (29)
If $v-v^{\prime}$ is $2s$-sparse and $J_{\ell}$ satisfies RIP, then by the RIP condition:
$(1-\delta_{2s})\|v-v^{\prime}\|_{2}^{2}\leq\|J_{\ell}(v-v^{\prime})\|_{2}^{2}\leq 4\epsilon^{2}$ (30)
Therefore:
$\|v-v^{\prime}\|_{2}\leq\frac{2\epsilon}{\sqrt{1-\delta_{2s}}}$ (31)
As $\epsilon\to 0$ (noiseless case), $v=v^{\prime}$, establishing uniqueness.
Conclusion (Sparsity): Under sparsity assumptions and RIP conditions—strong sufficient conditions that typically require careful measurement design—persona vectors are uniquely recoverable via $\ell_{1}$ minimization. RIP does not hold automatically for arbitrary Jacobians and must be verified or engineered through diverse prompt selection.
C.4 Proof of Condition: Multi-Environment Identification
Setup: Assume we observe the same persona $z$ across multiple environments $e\in\{1,\ldots,E\}$ where:
$h_{\ell}^{(e)}=g_{e}(x,z)+\epsilon^{(e)}$ (32)
The function $g_{e}$ may vary across environments (spurious correlations change) but the causal effect of $z$ on downstream outputs remains invariant.
Goal: Show that the invariant representation $z$ (and its associated direction) can be identified.
Step 1: Invariant causal mechanism.
Assume the causal graph:
[…]
where:
  • $z$: true persona (causal factor),
  • $x$: input prompt,
  • $e$: environment,
  • $h_{\ell}$: internal representation,
  • $o$: output.
The key assumption is invariance: the causal mechanism $h_{\ell}\to o$ is the same across environments, i.e., $F_{\ell\to L}$ does not depend on $e$.
Step 2: Invariant Risk Minimization (IRM).
The objective is to find representation $\Phi:\mathbb{R}^{d}\to\mathbb{R}^{k}$ and predictor $w:\mathbb{R}^{k}\to\mathbb{R}$ such that:
$\min_{\Phi,w}$ $\sum_{e=1}^{E}R^{e}(\Phi,w)$ (33)
s.t. $w\in\arg\min_{w^{\prime}}R^{e}(\Phi,w^{\prime})\quad\forall e$
where,
$R^{e}(\Phi,w)=\mathbb{E}_{(x,y)\sim\mathcal{P}^{e}}\left[\ell\big(w\cdot\Phi(h_{\ell}(x)),y\big)\right].$ (34)
Interpretation: Find a representation $\Phi(h_{\ell})=z$ such that the optimal predictor $w$ is the same across all environments. This filters out spurious correlations that vary with $e$.
Step 3: Identifiability under IRM.
Under the following idealized conditions:
  • Environments are sufficiently diverse: $\text{Cov}(x|e)$ varies across $e$
  • Causal mechanism is invariant: $p(o|z)$ is the same for all $e$
  • Sufficient environments: $E\geq k+1$ where $k=\dim(z)$
  • No unobserved confounders between environments and outcomes
The theorem (Ahuja et al., 2022; Von Kügelgen et al., 2021) states that the invariant risk minimization objective can recover $z$ up to an invertible transformation, though not necessarily uniquely.
Important caveat: The literature on invariant representation learning establishes conditions under which invariance constraints narrow the hypothesis class but full identifiability (unique recovery of $z$) requires additional assumptions beyond standard IRM formulations. The practical application of IRM to steering should be understood as identifying a stable, transferable representation rather than guaranteeing uniqueness. This is a sufficient condition in principle under strong assumptions, not a characterization of typical steering scenarios.
Step 4: Application to steering.
In practice:
  1. 1.
    Collect multi-environment data: Extract steering vectors from diverse prompt distributions, model checkpoints or instruction formats.
  2. 2.
    Learn invariant representation: Train $\Phi$ via IRM objective
  3. 3.
    Extract steering vectors: Compute $v_{i}=\nabla_{h_{\ell}}\Phi_{i}(h_{\ell})$ where $\Phi_{i}$ is the $i$-th component of $\Phi$
The resulting steering vectors $v_{i}$ capture invariant causal factors rather than spurious correlations.
Step 5: Theoretical guarantee.
Under mild conditions (sufficient diversity, invariance, identifiable causal structure), the IRM solution satisfies:
$\Phi(h_{\ell})=T(z)$ (35)
where $T$ is an invertible transformation. Since persona steering operates on $z$, the directions $v_{i}=\nabla_{h_{\ell}}\Phi_{i}$ recover the true causal factors.
Conclusion (Multi-environment): With diverse environments and invariant causal mechanisms—idealized sufficient conditions that narrow the hypothesis class—persona vectors corresponding to stable causal factors can be recovered up to invertible transformations. Full uniqueness requires additional assumptions beyond standard IRM formulations. This provides a principled approach to filtering spurious correlations but should be understood as identifying transferable representations rather than guaranteeing unique recovery.
C.5 Proof of Condition: Cross-Layer Consistency
Setup: Assume persona vectors exhibit consistent geometric relationships across layers:
$v_{\ell+1}=W_{\ell}v_{\ell}+\delta_{\ell}$ (36)
where $W_{\ell}\in\mathbb{R}^{d\times d}$ is the weight matrix connecting layers and $\|\delta_{\ell}\|$ is small.
Goal: Show that cross-layer constraints reduce the solution space and improve identifiability.
Step 1: Single-layer null space.
From proposition 1, at layer $\ell$:
$\mathcal{N}_{\ell}=\{v_{0}:J_{\ell}v_{0}=0\}$ (37)
with $\dim(\mathcal{N}_{\ell})=d-r_{\ell}$ where $r_{\ell}=\text{rank}(J_{\ell})$.
Step 2: Cross-layer propagation.
If $v_{\ell}\in\mathcal{N}_{\ell}$, does $v_{\ell+1}=W_{\ell}v_{\ell}\in\mathcal{N}_{\ell+1}$?
Generally, no. The null space changes across layers:
$v_{\ell}\in\mathcal{N}_{\ell}\not\Rightarrow W_{\ell}v_{\ell}\in\mathcal{N}_{\ell+1}$ (38)
Step 3: Intersection of constraints.
Consider steering vectors observed at multiple layers $\ell_{1},\ldots,\ell_{L}$. Each layer provides a constraint:
$J_{\ell_{i}}v_{\ell_{i}}=y_{i}$ (39)
If we additionally require consistency:
$v_{\ell_{i+1}}=W_{\ell_{i}}v_{\ell_{i}}+\delta_{\ell_{i}}$ (40)
Then the solution must satisfy:
$v_{\ell_{i}}\in\{v:J_{\ell_{i}}v=y_{i}\}\cap\{v:W_{\ell_{i}}v\approx v_{\ell_{i+1}}\}$ (41)
Step 4: Reduced null space (qualitative characterization).
The cross-layer constraints create an overdetermined system. The effective null space is:
$\mathcal{N}_{\text{eff}}=\bigcap_{i=1}^{L-1}\{v:W_{\ell_{i}}v\in\mathcal{N}_{\ell_{i+1}}\}$ (42)
Since each $W_{\ell_{i}}$ generically has full rank and maps null space vectors to non-null-space vectors, we expect:
$\dim(\mathcal{N}_{\text{eff}})\ll\dim(\mathcal{N}_{\ell})$
Intuition: Each additional layer imposes new independent constraints. If the propagation matrices $W_{\ell_{i}}$ are sufficiently "generic" (full rank with uncorrelated null space mappings), then each layer reduces the effective null-space dimension. For sufficiently many informative layers with uncorrelated constraint structures, the effective null space can shrink dramatically or even vanish.
Conclusion (Cross-layer): Cross-layer consistency constraints can substantially reduce null-space dimensionality by creating overdetermined systems. The extent of reduction depends on the specific geometric structure of propagation matrices and layer-wise Jacobians. While this approach does not guarantee complete identifiability in general, it provides a practical method for filtering spurious null-space components that lack geometric stability across layers.

### v2

@@ 6.3 Scale Invariance Analysis
The curves track closely with overlapping confidence bands, demonstrating that $v$ and $v+v_{\perp}$ produce statistically indistinguishable behavioral effects at all tested steering strengths. For Qwen2.5-3B, mean differences remain below $0.022$ across all $\alpha$ values (less than $2.5\%$ deviation in formality scores); for Llama-3.1-8B, differences remain below $0.023$ ($2.5\%$ deviation). This proves that non-identifiability is a structural property: any vector in the equivalence class $v+\ker(J)$ produces identical observations regardless of scaling.
6.4 Robustness of Non-Identifiability Under Distribution Shift
We test whether equivalence classes persist across distributional shifts or whether prompt diversity alone induces identifiability. We evaluate orthogonal perturbation equivalence across four distinct prompt environments: in-distribution (standard trait templates), topic shift (medical, legal, creative writing contexts), genre shift (emails, tweets, academic writing), and safety-style (refusal patterns, ethical dilemmas). For each environment, we extract steering vectors and evaluate whether $v\equiv v+v_{\perp}$ equivalence holds when testing on that environment’s prompts.
Table 2 and Figure 4 present Cohen’s $d$ effect sizes across all trait-environment combinations for both models. Across 24 conditions, we observe effect sizes ranging from $d=0.21$ to $d=0.63$ (mean $d=0.35$). Effect sizes are predominantly small ($d<0.5$ for 88% of conditions; mean $d=0.35$), indicating that $v$ and $v+v_{\perp}$ remain largely observationally equivalent across diverse extraction contexts.
Table 2: Multi-environment validation. Cohen’s $d$ effect sizes across extraction contexts. ID = in-distribution; Top = topic shift; Gen = genre shift; Saf = safety-style prompts.
Model Trait ID Top Gen Saf
Qwen-3B Formality 0.44 0.25 0.29 0.21
Politeness 0.22 0.24 0.52 0.41
Humor 0.42 0.51 0.24 0.51
Llama-8B Formality 0.30 0.54 0.30 0.21
Politeness 0.37 0.27 0.51 0.63
Humor 0.29 0.56 0.30 0.40
Figure 4: Multi-environment validation results. Environment-dependent effect sizes and cross-environment matrices show small-to-moderate Cohen’s $d$ values across traits and prompt contexts, confirming robustness of observational equivalence under distributional shift.
Formality shows most consistency (mean $d=0.32$), with particularly strong equivalence in safety-style contexts ($d=0.21$ for both models). Politeness exhibits higher variability ($d$ reaching 0.52–0.63 in genre-shift and safety contexts) but remains within the small-effect range. Humor demonstrates intermediate stability (mean $d=0.39$), with genre-shift producing notably lower effects ($d=0.24$ for Qwen, $d=0.30$ for Llama). Figure 4 visualizes these patterns through environment-dependent effect sizes (left panels) and cross-environment transfer matrices (right panels), showing that equivalence remains stable even when vectors are extracted in one environment and tested in another.
Implications. The persistence of small effect sizes across diverse prompt distributions demonstrates that non-identifiability is not an artifact of in-distribution prompts but a fundamental geometric property. Critically, this shows that naive distributional diversity without explicit invariance constraints or structural assumptions is insufficient to recover identifiable representations.
@@ 7 Conclusion
We have established that steering vectors are fundamentally non-identifiable under standard observational regimes, with infinitely many distinct directions producing equivalent behavioral effects. This non-identifiability persists across models, traits, and prompt distributions, revealing a fundamental tension: methods optimized for behavioral control may not yield interpretable representations. Our work clarifies when steering provides reliable heuristic modification versus principled causal understanding, highlighting that interpretability claims require validation beyond input-output equivalence alone.
@@ 8 Limitations
Our empirical validation covers two models at mid-network layers across three semantic traits (formality, politeness, humor). While our theoretical results apply generally, the empirical magnitude of non-identifiability may vary across model families, scales, architectures and layer positions. Our semantic evaluation uses lexical heuristics rather than full distributional metrics or human judgments, providing a conservative test of observational equivalence.
Our multi-environment experiments demonstrate that equivalence persists across diverse prompt distributions, but we do not test whether explicit structural interventions such as independence constraints (ICA), sparsity regularization, or invariance objectives (IRM) can recover identifiable representations. While theoretical frameworks suggest these approaches might enable identifiability, empirical validation across multiple models and experimental designs remains an important direction for future work. Additionally, we do not directly measure null-space dimensionality or test whether null-space versus row-space components exhibit differential robustness under distribution shift.
@@ 9 Future Work
Our results establish the necessity of validation beyond input-output behavior. Important directions include testing whether structural constraints can recover identifiable representations, analyzing how non-identifiability varies across network layers and architectural components, and characterizing how null-space dimensionality scales with model size.
Investigating whether alternative steering methods exhibit superior robustness under adversarial attacks or distribution shift would clarify the practical consequences of non-identifiability. Human evaluation would provide stronger validation than our lexical heuristics, and direct measurement of null-space dimensionality would quantify our theoretical predictions.
10 Acknowledgments
@@
Appendix A Logit-Level Equivalence Test
To complement our semantic evaluations, we perform a direct logit-level equivalence test. For each model and trait, we compare the next-token logit vectors produced by steering with $v$ and with $v+v_{\perp}$, where $v_{\perp}$ is a norm-matched direction orthogonal to $v$. We measure the $\ell_{2}$ distance between logit vectors and normalize by the corresponding distance induced by a norm-matched random direction:
$R=\frac{\|\ell(v)-\ell(v+v_{\perp})\|_{2}}{\|\ell(v)-\ell(v_{\text{rand}})\|_{2}}$ (10)
Values $R<1$ indicate that orthogonal perturbations induce smaller logit deviations than random directions. In all tested models and traits, this condition holds. Ratios range from $0.49$ (Llama, formality) to $0.90$ (Qwen, humor), indicating that orthogonal perturbations systematically induce smaller logit shifts than random directions. The next-token predictions remain highly stable: token agreement between $v$ and $v+v_{\perp}$ ranges from $81\%$ to $96\%$, and top-10 token overlap ranges from $86\%$ to $96\%$.
These results provide direct evidence that equivalence is not an artifact of semantic measurement. While orthogonal perturbations induce some logit deviation, they systematically preserve output distributions more faithfully than random directions, consistent with Proposition 1’s prediction of high-dimensional equivalence classes. Table 3 reports mean $\ell_{2}$ logit distances, normalized ratios, token agreement, and top-$k$ overlap for all models and traits.
Table 3: Logit-level equivalence statistics.
Model Trait $R$ (v_perp / random) Token Agreement Top-10 Overlap
Qwen-3B Humor 0.90 81% 86%
Qwen-3B Formality 0.76 94% 96%
Qwen-3B Politeness 0.75 93% 96%
Llama-8B Humor 0.64 96% 90%
Llama-8B Formality 0.49 93% 95%
Llama-8B Politeness 0.65 91% 94%
Appendix B Prompt Construction Details
B.1 Contrastive Prompt Pairs
We construct contrastive prompt pairs for each semantic trait using template-based generation. Table 4 provides concrete examples for all three traits used in the empirical validation.
Trait Positive ($x^{+}$) Negative ($x^{-}$)
Formality Draft a professional email summarizing quarterly performance. Text a friend about how the quarter went.
Politeness Respond diplomatically to criticism from a colleague. Reply bluntly to someone who criticized you.
Humor Tell a lighthearted story about an embarrassing mistake. Provide a factual account of an embarrassing mistake.
Table 4: Example contrastive prompt pairs for each semantic trait.
B.2 Steering Extraction and Evaluation Pipeline
Figure 5: Steering extraction and evaluation pipeline.
@@ B.2 Steering Extraction and Evaluation Pipeline
Appendix C Detailed Proof of Proposition 1
We provide proofs for establishing non-identifiability via two complementary arguments: null-space ambiguity and reparameterization symmetry.
C.1 Null-Space Ambiguity (Constructive Proof)
@@
$o(x,v^{\prime},\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}=o(x,0,0)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha).$ (18)
@@
C.2 Reparameterization Symmetry (Existence Proof)
@@
$h^{\prime}_{\ell+1}=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)=h_{\ell+1}.$ (25)
@@
$\tilde{h}^{\prime}_{\ell}=h^{\prime}_{\ell}+\alpha v^{\prime}=Ah_{\ell}+\alpha Av=A(h_{\ell}+\alpha v)=A\tilde{h}_{\ell}$. (28)
$\tilde{h}^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}(h^{\prime}_{\ell}+\alpha v^{\prime})+b^{\prime}_{\ell+1}\right)$ (29)